solve the equations
-143=7x-3(2x+13)
5x+3/9=-3

Answer by TimothyLamb(4379)

-143 = 7x - 3(2x + 13)
-143 = 7x - 6x - 13*3
7x - 6x = -143 + 39
x = -104
---
5x + 3/9 = -3
5x = -3 - (3/9)
5x = -(3*9/9) - (3/9)
5x = -(27/9) - (3/9)
5x = -(30/9)
x = -(30/9)/5
x = -( 3*2*5/(3*3) )/5
x = -2/3
x = -0.66666666666
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---